Identify and explain the kinematics in rotational motion. Compare the kinematics in rotational


motion to the equivalent translational motion

1
Expert's answer
2021-12-20T10:31:31-0500

The kinematics in rotational motion is the field of mechanics that ties together the variables that define the rotational motion: angular displacement "\phi", angular velocity "\omega", angular acceleration "\alpha":


"\alpha=\frac{\omega_f^2-\omega_i^2}{2\phi},\space a=\frac{v_f^2-v_i^2}{2d}.\\\space\\
\omega=\sqrt{2\alpha \phi},\space v=\sqrt{2ad}.\\\space\\
\phi=\frac 12 \alpha t^2,\space d=\frac12at^2."